# On the Curvature of Regge Metrics

### v1

@@ Appendix B Appendix: Relationship with Existing Formulas
$\frac{1}{2}\Theta_{p}$ $\sum_{k,l}\left\langle\tau_{n}^{T}\otimes{\tau^{n-1}}^{T}-\tau_{n-1}^{T}\otimes{\tau^{n}}^{T},\phi^{i}_{~jkl}\tau_{i}^{T}\otimes{\tau^{j}}^{T}\right\rangle i_{p}^{*}\star({\tau^{k}}^{T}\wedge{\tau^{l}}^{T})$
@@
$=\frac{1}{2}\Theta_{p}\left\langle\tau_{n}^{T}\otimes{\tau^{n-1}}^{T}-\tau_{n-1}^{T}\otimes{\tau^{n}}^{T},[\phi^{i}_{~j,n-1,n}-\phi^{i}_{~j,n,n-1}]\tau_{i}^{T}\otimes{\tau^{j}}^{T}\right\rangle\omega_{p}$
$=\Theta_{p}\left\langle\tau_{n}\otimes{\tau^{n-1}}^{T}-\tau_{n-1}\otimes{\tau^{n}}^{T},\phi^{i}_{~j,n-1,n}\tau_{i}^{T}\otimes{\tau^{j}}^{T}\right\rangle\omega_{p}$
$=\Theta_{p}(\delta_{n}^{j}\delta_{i}^{n-1}-\delta_{n-1}^{j}\delta_{i}^{n})\phi^{i}_{~j,n-1,n}\omega_{p}$
$=\Theta_{p}(\phi^{n-1}_{~n,n-1,n}-\phi^{n}_{~n-1,n-1,n})\omega_{p}$
$=2\Theta_{p}\phi^{n-1}_{~n,n-1,n}\omega_{p}=2\Theta_{p}A(\tau_{n-1}^{T},\tau_{n}^{T},\tau_{n}^{T},\tau_{n-1}^{T})\omega_{p}=2\Theta_{p}A_{\hat{\mu}\hat{\nu}\hat{\nu}\hat{\mu}}\omega_{p}.$
@@
Let $\hat{\phi}\in\mathcal{A}(f)$ and let $A$ be defined as in Proposition 1. Then
@@
The only term which is not straightforward from Proposition 1 and the definition of $\widetilde{\mathcal{R}\omega}$ in [8, p. 11] is the term involving the jump in second fundamental form across codimension-1 interfaces.
@@
$\int_{\mathring{e}}\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{e}^{*}\hat{\phi}^{T}\right\rangle-\left\langle\hat{\mathrm{I\!I}}_{e}^{T^{\prime}}\wedge i_{e}^{*}\hat{\phi}^{T^{\prime}}\right\rangle,$

### v2

@@ Appendix B Appendix: Relationship with Existing Formulas
$\frac{1}{2}\Theta_{p}$ $\sum_{i,j,k,l}\left\langle\tau_{n}^{T}\otimes{\tau^{n-1}}^{T}-\tau_{n-1}^{T}\otimes{\tau^{n}}^{T},\phi^{i}_{~jkl}\tau_{i}^{T}\otimes{\tau^{j}}^{T}\right\rangle i_{\mathring{p}}^{*}\star({\tau^{k}}^{T}\wedge{\tau^{l}}^{T})$
@@
$=\frac{1}{2}\Theta_{p}\sum_{i,j}\left\langle\tau_{n}^{T}\otimes{\tau^{n-1}}^{T}-\tau_{n-1}^{T}\otimes{\tau^{n}}^{T},[\phi^{i}_{~j,n,n-1}-\phi^{i}_{~j,n-1,n}]\tau_{i}^{T}\otimes{\tau^{j}}^{T}\right\rangle\omega_{p}$
$=\Theta_{p}\sum_{i,j}\left\langle\tau_{n}\otimes{\tau^{n-1}}^{T}-\tau_{n-1}\otimes{\tau^{n}}^{T},\phi^{i}_{~j,n,n-1}\tau_{i}^{T}\otimes{\tau^{j}}^{T}\right\rangle\omega_{p}$
$=\Theta_{p}\sum_{i,j}(\delta_{n}^{j}\delta_{i}^{n-1}-\delta_{n-1}^{j}\delta_{i}^{n})\phi^{i}_{~j,n,n-1}\omega_{p}$
$=\Theta_{p}(\phi^{n-1}_{~n,n,n-1}-\phi^{n}_{~n-1,n,n-1})\omega_{p}$
$=2\Theta_{p}\phi^{n-1}_{~n,n,n-1}\omega_{p}=2\Theta_{p}A|_{T,e}(\tau_{n}^{T},\tau_{n-1}^{T},\tau_{n-1}^{T},\tau_{n}^{T})\omega_{p}=2\Theta_{p}A_{\hat{\mu}\hat{\nu}\hat{\nu}\hat{\mu}}\omega_{p}.$
@@
Let $\hat{\phi}\in\mathcal{A}(f,\mathcal{T},M)$ and let $A$ be defined as in Proposition 16. Then, using notation from this paper on the left side and notation from [13] on the right side,
@@
The only term which is not straightforward from Proposition 16 and the definition of $\widetilde{\mathcal{R}\omega}$ in [13, p. 11] is the term involving the jump in second fundamental form across codimension-1 interfaces.
@@
$\int_{\mathring{e}}\left\langle\hat{\mathrm{I\!I}}_{e}^{T}\wedge i_{\mathring{e}}^{*}\hat{\phi}^{T}\right\rangle-\left\langle\hat{\mathrm{I\!I}}_{e}^{T^{\prime}}\wedge i_{\mathring{e}}^{*}\hat{\phi}^{T^{\prime}}\right\rangle,$